OpenStudy (anonymous):

Simplify the expression by using a double-angle formula. cos^2(theta/3) - sin^2(theta/3)

OpenStudy (anonymous):

use cos^2 x -sin^2x =cos2x

OpenStudy (anonymous):

How this is new for me

OpenStudy (anonymous):

I need help please

OpenStudy (anonymous):

do u know cos(a+b)=cosa cosb -sina sinb now put a=x and b=x so we have cos(x+x)=cosx cosx -sinx sinx or cos2x =(cosx)^2 -(sinx)^2 or cos2x = cos^2 x -sin^2x

OpenStudy (anonymous):

i will try it one moment

OpenStudy (anonymous):

so the answer will be cos(2.theta/3) right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Thanks matricked